# Enhance Lifelong Model Editing with Continuous Data-Adapter Association

### v1

@@ 3.1.3 Datasets
Base FT-L ROME Defer MELO GRACE ELDER
ZsRE GPT2-XL Reliability 0.00 48.23 48.40 52.95 70.81 96.80 97.47
Generalization 0.00 47.61 47.20 44.83 66.41 0.00 96.08
LLaMA2-7B Reliability 0.25 32.21 77.60 53.82 64.57 89.48 93.96
Generalization 0.37 28.96 74.53 51.32 42.93 0.46 90.21
CounterFact GPT2-XL Reliability 0.00 55.10 69.00 58.80 65.80 88.90 94.65
Generalization 0.00 44.06 68.74 44.43 49.20 76.05 91.26
LLaMA2-7B Reliability 0.40 66.26 79.37 68.82 51.39 77.70 95.07
Generalization 0.24 44.65 80.30 57.75 35.71 63.35 90.79
@@
To evaluate the test retention of post-edit LLMs on general tasks, we expand the evaluation task data used in previous studies (Hartvigsen et al., 2024; Yu et al., 2024) for a more comprehensive assessment of the post-edit LLMs’ general abilities. This expansion is crucial because recent studies show that current model editing methods can degrade LLM performance on downstream tasks (Gu et al., 2024; Gupta et al., 2024; Yang et al., 2024), underscoring the importance of assessing their side effects with diverse datasets. Specifically, we employ a benchmark from (Gu et al., 2024), including eight diverse tasks: Reasoning on GSM8K (Cobbe et al., 2021), Natural Language Inference on RTE(Dagan et al., 2005), Open-domain QA on Natural Question(Kwiatkowski et al., 2019), Closed-domain QA on BoolQ(Clark et al., 2019), Dialogue on MuTual(Cui et al., 2020), Summarization on SAMSum (Cui et al., 2020), Named Entity Recognition on CoNLL03(Sang and De Meulder, 2003), and Sentiment Analysis on SST2(Socher et al., 2013). Their respective metrics are in Appendix B.2.
3.1.4 Implement Details
We use two LLMs as base models: LLaMA2 (7B) (Touvron et al., 2023) and GPT2-XL (1.5B)(Radford et al., 2019). To evaluate our baselines, we use their original implementations and adapt them to our datasets and base models. For our proposed ELDER across all settings, the rank of LoRAs is set to 8, and the number of layers that apply mixture-of-LoRA is set to 6. The number of LoRAs per layer is set to 4, $k$ is set to 2, and $\epsilon$ is set to 12. $\lambda$ is set to $21e-$ 2. More training details are available in Appendix A.
3.2 Experimental Results
3.2.1 Main Results
We compare the lifelong model editing performance of our proposed ELDER with recently advanced baselines in Table 1. We observe that ELDER consistently outperforms these baselines after long sequences of edits across all datasets and base models. FT-L and ROME, which are not designed for sequential editing, tend to forget previous edits. GRACE, which relies on discrete data-adapter mapping, is a strong baseline for reliably remembering previous edits but lacks robustness when handling semantically equivalent inputs, leading to poor generalization scores, especially on ZsRE, as noted in previous works (Tian et al., 2024; Lin et al., 2024). In contrast, ELDER excels in maintaining previous edits and robustly generalizes to their rephrases, demonstrating its superiority.
@@ 3.2.1 Main Results
Table 2 reports the test retention on general tasks after sequential edits by each method. Detailed results for each task are provided in Appendix C. ELDER effectively retains the LLM general abilities after lifelong editing, preserving the practical value of post-edit models. The results demonstrate that our deferral mechanism successfully identifies task inputs using edit-related information from LoRA allocations. On the other hand, FT-L and ROME significantly degrade LLM performance due to repeated modifications of model parameters, consistent with previous studies (Gu et al., 2024; Gupta et al., 2024).
3.2.2 Editing Efficiency and Scalability
Efficiency and scalability are critical for lifelong model editing methods, as they facilitate the management of multiple sequential edits and accommodate the continual increase in edit sequence length. We compare the editing efficiency of ELDER against GRACE in Table 3, showing the number of learnable parameters and average editing time after 1000 sequential edits from the ZsRE dataset. The results clearly indicate that ELDER is more efficient than the baseline method. ELDER’s time efficiency stems from its end-to-end design, which avoids the need to maintain clusters and search for discrete mappings.
GPT2 LLaMA2
@@ 3.2.2 Editing Efficiency and Scalability
We also examine how ELDER scales with different parameter budgets. Figure 3 presents the results of varying $L$, which represents the number of layers utilizing the mixture-of-LoRA module, initially set to $6$. As $L$ increases, ELDER becomes more stable and demonstrates better scalability, benefiting from the increased number of learnable parameters.
@@
3.2.3 Ablation Experiments
We conduct ablation experiments to evaluate the effect of our proposed guided loss. Two alternatives are considered: removing the guided loss during training or replacing it with the load balancing loss proposed by Fedus et al. (2021). The latter improves the routing scheme by calculating the average LoRA usage load per training batch and optimizes it to be uniformly distributed. The results, shown in Table4, demonstrate that the guided loss achieves superior results and enhances the training of the mixture-of-LoRAs for lifelong model editing. In contrast, directly uniforming usage load per batch degrades the performance, since the small batch size in lifelong editing leads to the disparity between batch measurement and real usage loads, as described in Section 2.3.
@@ 3.2.3 Ablation Experiments
Table 4: Editing generalization while altering the auxiliary loss. w/o guide denotes removing guided loss from training. w balancing means balancing load with the loss proposed by Fedus et al. (2021) instead.
3.2.4 Model Analysis
We further assess the behavior of ELDER by visualizing the LoRA allocation codes generated by the mixture-of-LoRA modules. For our experiments, we use ZsRE and LLaMA2-7B. We randomly sample five groups of semantically equivalent inputs, including edits and their corresponding rephrases, along with twenty samples from the Natural Question dataset, chosen as task inputs unrelated to the edited knowledge. The allocation codes for these inputs are recorded and visualized using t-SNE, as shown in Figure 4. We observe that semantically equivalent inputs receive similar, though not identical, LoRA allocations. This similarity suggests that ELDER effectively captures the semantic associations between edits and their rephrases, showing robustness to minor input variations. Additionally, task inputs are distinctly separated from all edits, supporting the use of LoRA allocation codes for discrimination in our deferral mechanism.
@@ 3.2.4 Model Analysis
4 Related Works
Model Editing
Model editing aims to accurately and efficiently modify knowledge in deep neural networks(Sinitsin et al., 2019), especially language models(Yao et al., 2023). Meta-learning-based methods MEND(Mitchell et al., 2021) and KE(De Cao et al., 2021) train an extra hypernetwork to learn changes in the base model. KN(Dai et al., 2022a) attributes the neuron that embodies the knowledge and updates these neurons. ROME(Meng et al., 2022) locates the edit area by casual analysis and modifies the entire weight matrix. SERAC(Mitchell et al., 2022) preserves model parameters and introduces a counterfactual model to process updated knowledge.
However, the above methods only consider static edits, i.e., modifying the model a single time. Huang et al. (2022) proposed a sequential editing approach by adding one neuron for each edit sample. Nevertheless, this method relies on large sets of unrelated inputs and is slow due to the need to train neurons for each edit (Yao et al., 2023). Besides, it is architecture-specific and lacks a clear extension to GPT-based models. Most similar works to ours are GRACE (Hartvigsen et al., 2024) and MELO(Yu et al., 2024), which avoid interfering with successive edits by discretely mapping different adapters to edit samples.
Mixture-of-Experts (MoE)
Jacobs et al. (1991); Jordan and Jacobs (1994) introduced MoE models to compute different examples with independent expert modules. Shazeer et al. (2016) extended this concept to large-scale language models with LSTMs. Advances such as GShard (Fedus et al., 2021), Switch Transformer (Fedus et al., 2021), and BASE Layer Lewis et al. (2021) have refined routing input tokens to experts. Hash Layer (Roller et al., 2021) uses a pre-defined token-level hash table for token-to-expert assignment. StableMoE (Dai et al., 2022b) addresses routing fluctuation by training the router network first and then freezing it for further model training. In contrast, our approach guides the router network to learn a pre-set sample-to-adapter assignment, effectively handling unseen but equivalent inputs.
LoRA Combinations
LoRA (Hu et al., 2021) uses low-rank matrices to update the LLMs, and has become a popular parameter-efficient fine-tuning (PEFT) method. Recently, researchers have been using combinations of multiple LoRAs for further benefits. Huang et al. (2023) proposed composing existing LoRA modules for generalization on new tasks. Zadouri et al. (2023) combined LoRA with token-level mixture-of-experts (Jacobs et al., 1991) and designed a new PEFT module. Dou et al. (2023) split LoRA experts into two separate groups to maintain world knowledge of LLMs during instruction tuning. While these works apply different experts at the token level, we focus on routing edit data samples to different adapters. To the best of our knowledge, no study has focused on improving current lifelong model editing methods with a multi-LoRA architecture.
5 Conclusion
@@
In conclusion, this paper presents a novel method for lifelong model editing utilizing mixture-of-LoRAs. It effectively enhances editing robustness by associating LoRA allocations with edit semantics smoothly. Extensive experiments have shown that this method significantly enhances editing performance with efficiency and scalability while simultaneously preserving LLMs’ original performance on general tasks.
Limitations
This paper uses the datasets referenced in previous work and focuses on evaluating the generalization and robustness of model editing methods through rephrased edits. However, it remains underexplored how models perform after lifelong editing on more complex data containing edited knowledge. This presents an open area for future research.
Appendix A Training Details
@@
We provide more training details for our experiments. All methods are trained on a 48GB NVIDIA A40, and editing efficiency is evaluated on the same device. We train ELDER with the Adam optimizer (Kingma and Ba, 2014). For GPT2-XL, we edit the fully-connected component from the thirteen-fifth to fortieth layers (from $transformer.h.36.mlp.c\_fc$ to $transformer.h.40.mlp.c\_fc$). For LLaMA2-7B, we edit the fully connected component from the twentieth to the twenty-fifth layers (from $model.layers.21.mlp.down\_proj$ to $model.layers.26.mlp.down\_proj$). On all experiment settings, the learning rate of training ELDER is set to $41e-$ 4, the batch size is set to $4$, and we train for $50$ iterations.
Appendix B Additional Dataset Descriptions
B.1 Preprocessing Model Editing Dataset
@@
We choose ZsRE and CounterFact for experiments and utilize them following the methodologies outlined in (Hartvigsen et al., 2024) and (Yu et al., 2024). We use the same data to evaluate all baselines and ELDER. For ZsRE, we follow the previous data split and evaluate all models on the validation set. For CounterFact, we use the original data file. For both datasets, we extract the first 1000 edits with at least one rephrased data sample for the main experiments. We split the edits and their rephrasings into two groups, using the first group to edit the model sequentially and the latter group to evaluate the editing generalization.
B.2 Metrics of Test Retention Dataset
@@
  • Reasoning on GSM8K (Cobbe et al., 2021). The results are measured by the solve rate.
  • Natural Language Inference on RTE(Dagan et al., 2005). The results are measured by two-way classification.
  • Open-domain QA on Natural Question(Kwiatkowski et al., 2019). The results are measured by accuracy with the reference answer.
  • Closed-domain QA on BoolQ(Clark et al., 2019). The results are measured by exact match.
  • Dialogue on MuTual(Cui et al., 2020). The results are measured by selecting one best-matched response from four available choices.
  • Summarization on SAMSum (Cui et al., 2020). The results are measured by the average of ROUGE-1, ROUGE-2, and ROGUE-L.
  • Named Entity Recognition on CoNLL03(Sang and De Meulder, 2003). The results are measured by entity-level F1-score.
  • Sentiment Analysis on SST2(Socher et al., 2013). Results are measured by the accuracy of two-way classification.
Appendix C Detailed Results of Test Retention on General Tasks
The detailed results of the pre-edit and post-edit models on downstream tasks under each experimental setting are shown in Table 5, Table 6, Table 7 and Table 8.. Each task is described in Section 3.1.3. The last column in each table stands for the standard test retention we report in Table 2.
@@ Appendix C Detailed Results of Test Retention on General Tasks
Table 5: Results on general task of LLaMA-7B after editing with ZsRE

### v2

@@ Introduction
Base FT-L LoRA ROME WilKE SERAC T-Patcher MELO GRACE ELDER
ZsRE GPT2-XL Reliability 0.00 48.23 30.22 48.40 53.32 96.09 77.29 70.81 96.80 97.47
Generalization 0.00 47.61 19.39 47.20 47.10 53.03 67.74 66.41 0.00 96.08
LLaMA2-7B Reliability 0.25 32.21 10.70 77.60 61.84 83.14 62.94 64.57 89.48 93.96
Generalization 0.37 28.96 7.31 74.53 51.29 60.38 48.37 42.93 0.46 90.21
CounterFact GPT2-XL Reliability 0.00 55.10 37.21 69.00 68.19 100.00 91.36 65.80 88.90 94.65
Generalization 0.00 44.06 34.55 68.74 53.35 79.49 80.31 49.20 76.05 91.26
LLaMA2-7B Reliability 0.40 66.26 10.16 79.37 74.78 82.67 88.93 51.39 77.70 95.07
Generalization 0.24 44.65 5.21 80.30 55.83 71.83 77.75 35.71 63.35 90.79
@@
To evaluate the test retention of post-edit LLMs on general tasks, we expand the evaluation task data used in previous studies (Hartvigsen et al. 2024; Yu et al. 2024) for a more comprehensive assessment of the post-edit LLMs’ general abilities. This expansion is crucial because recent studies show that current model editing methods can degrade LLM performance on downstream tasks (Gu et al. 2024; Gupta, Rao, and Anumanchipalli 2024; Yang et al. 2024b), underscoring the importance of assessing their side effects with diverse datasets. Specifically, we employ a benchmark from (Gu et al. 2024), including eight diverse tasks: Reasoning on GSM8K (Cobbe et al. 2021), Natural Language Inference on RTE(Dagan, Glickman, and Magnini 2005), Open-domain QA on Natural Question(Kwiatkowski et al. 2019), Closed-domain QA on BoolQ(Clark et al. 2019), Dialogue on MuTual(Cui et al. 2020), Summarization on SAMSum (Cui et al. 2020), Named Entity Recognition on CoNLL03(Sang and De Meulder 2003), and Sentiment Analysis on SST2(Socher et al. 2013). Their respective metrics are in the technical appendix.
Implement Details
We use two LLMs as base models: LLaMA2 (7B) (Touvron et al. 2023) and GPT2-XL (1.5B)(Radford et al. 2019). To evaluate our baselines, we use their original implementations and adapt them to our datasets and base models. For our proposed ELDER across all settings, the rank of LoRAs is set to 8, and the number of layers that apply mixture-of-LoRA is set to 6. The number of LoRAs per layer is set to 4, $k$ is set to 2, and $\epsilon$ is set to 12. $\lambda$ is set to $21e-$ 2. More details of training and hyperparameter tuning are available in the technical appendix.
Experimental Results
Main Results
We compare the lifelong model editing performance of our proposed ELDER with recently advanced baselines, as shown in Table 1. Initially, the base models show low performance, indicating that the edit knowledge is not inherent to the model. After long sequences of edits, we observe that ELDER consistently outperforms these baselines, across all datasets and base models. FT-L and ROME, which are not designed for sequential editing, tend to forget previous edits. GRACE, which relies on discrete data-adapter mapping, is a strong baseline for reliably remembering previous edits but lacks robustness when handling semantically equivalent inputs, leading to poor generalization scores, especially on ZsRE, as noted in previous works (Tian et al. 2024; Lin et al. 2024). SERAC and T-Patcher also exhibit good lifelong editing reliability, but fall short in generalization. In contrast, ELDER excels in maintaining previous edits and robustly generalizes to their rephrases, demonstrating its superiority.
@@
Table 2 reports the test retention on general tasks after sequential edits by each method. Detailed results for each task are provided in the technical appendix. ELDER effectively retains the LLM general abilities after lifelong editing, preserving the practical value of post-edit models. The results demonstrate that our deferral mechanism successfully identifies task inputs using edit-related information from LoRA allocations. On the other hand, FT-L, LoRA, ROME, WilKE and T-Patcher significantly degrade LLM performance due to repeated modifications of model parameters, consistent with previous studies (Gu et al. 2024; Gupta, Rao, and Anumanchipalli 2024).
Editing Efficiency and Scalability
Efficiency and scalability are critical for lifelong model editing methods, as they enable the model to manage multiple sequential edits with acceptable cost, when the edit sequence length continually increases. We compare the editing efficiency of ELDER against GRACE and SERAC in Table 3. Both methods exhibit strong performance in lifelong model editing and the retention of general abilities. We show the number of extra parameters and average editing time after 1000 sequential edits from the ZsRE dataset. The results clearly indicate that ELDER is more efficient than the baseline method. ELDER’s time efficiency stems from its end-to-end design, which avoids the need to search for discrete mappings compared with GRACE. The editing speed of SERAC is not reported, because it requires additional training of the scope classifier and counterfactual model for each new edit. Moreover, we show the inference efficiency of ELDER in the technical appendix.
GPT2-XL LLaMA2-7B
@@
We also examine how ELDER scales with different parameter budgets. Figure 3 presents the results of varying $L$, which represents the number of layers utilizing the mixture-of-LoRA module, initially set to $6$. As $L$ increases, ELDER becomes more stable and exhibits better scalability, benefiting from the increased number of learnable parameters.
@@
Ablation Experiments
We conduct ablation experiments to evaluate the effect of our proposed guided loss. Two alternatives are considered: removing the guided loss during training or replacing it with the load balancing loss proposed by Fedus, Zoph, and Shazeer (2021). The latter improves the routing scheme by calculating the average LoRA usage load per training batch and optimizes it to be uniformly distributed. The results, shown in Table4, demonstrate that the guided loss achieves superior results and enhances the training of the mixture-of-LoRAs for lifelong model editing. In contrast, directly uniforming usage load per batch degrades the performance, since the small batch size in lifelong editing leads to the disparity between batch measurement and real usage loads, as described in the Method SectionGuided Loss.
@@
Table 4: Editing generalization while altering the auxiliary loss. w/o guide denotes removing guided loss from training. w balancing means balancing load with the loss proposed by Fedus, Zoph, and Shazeer (2021) instead.
Model Analysis
We further assess the behavior of ELDER by visualizing the LoRA allocation codes generated by the mixture-of-LoRA modules. We use ZsRE and LLaMA2-7B for the analysis. We randomly sample five groups of semantically equivalent inputs, including edits and their corresponding rephrases, along with twenty samples from the Natural Question dataset, chosen as task inputs unrelated to the edited knowledge. The allocation codes for these inputs are recorded and visualized using t-SNE, as shown in Figure 4. We observe that semantically equivalent inputs receive similar, though not identical, LoRA allocations. This similarity suggests that ELDER effectively captures the semantic associations between edits and their rephrases, showing robustness to minor input variations. Additionally, task inputs are separated from all edits, supporting the use of allocation codes for discrimination in our deferral mechanism.
@@
Conclusion
@@
In conclusion, this paper presents ELDER, a novel approach for lifelong model editing using a mixture-of-LoRA structure. The training process is guided by a novel loss function to align the LoRA allocation with edit knowledge, and a deferral mechanism is integrated to retain general abilities after editing. ELDER improves editing generalization by effectively assigning similar LoRA allocations to semantically equivalent inputs during inference. Extensive experiments demonstrate that this method significantly enhances editing performance with remarkable efficiency and scalability, while also preserving the original performance of LLMs on general tasks.
Acknowledgements
We would like to thank Rui-chen Zheng for the discussion on optimizing adapter allocation. This research is supported by Artificial Intelligence-National Science and Technology Major Project 2023ZD0121200 and National Natural Science Foundation of China under Grant No.62222212, No.62376033, No.62372006.
Appendix
Training Details
@@ Acknowledgements
We provide more training details for our experiments here. All methods are trained on a single 48GB NVIDIA A40, and editing efficiency is evaluated on the same device. We train ELDER with the Adam optimizer (Kingma and Ba 2014).
The default hyperparameter setting we use are as follows. For GPT2-XL, we edit the fully-connected component from the thirteen-fifth to fortieth layers (from $transformer.h.36.mlp.c\_fc$ to $transformer.h.40.mlp.c\_fc$). For LLaMA2-7B, we edit the fully-connected component from the twentieth to the twenty-fifth layers (from $model.layers.21.mlp.down\_proj$ to $model.layers.26.mlp.down\_proj$). On all experiment settings, the learning rate of training ELDER is set to $41e-$ 4, the batch size is set to $4$, and we train for $50$ iterations.
For hyperparameter tuning, we list the tried values of several hyperparameters in ELDER, in the Table 5. During the development of this paper, we used Generalization as the criterion when deciding the better $k$, $N$, and $\lambda$, and used Test Retention on General Tasks when altering $\epsilon$.
Hyperparameters Values
k {1, 2, 3, 4, 5}
N {2, 4, 8, 16}
$\lambda$0.001, 0.01, 0.1, 1, 10}
$\epsilon$0, 6, 12, 18, 24}
Table 5: Hyperparameter Values
Additional Dataset Descriptions
Preprocessing Model Editing Dataset
@@
We choose ZsRE and CounterFact for experiments and utilize them following the methodologies outlined in (Hartvigsen et al. 2024) and (Yu et al. 2024). We use the same data to evaluate all baselines and ELDER. For ZsRE, we follow the previous data split and evaluate all models on the validation set. For CounterFact, we use the original data file. For both datasets, we extract the first 1000 edits with at least one rephrased data sample for the main experiments. We split the edits and their rephrasings into two groups, using the first group to edit the model sequentially and the latter group to evaluate the editing generalization.
Metrics of Test Retention Dataset
@@
  • Reasoning on GSM8K (Cobbe et al. 2021). The results are measured by the solve rate.
  • Natural Language Inference on RTE(Dagan, Glickman, and Magnini 2005). The results are measured by two-way classification.
  • Open-domain QA on Natural Question(Kwiatkowski et al. 2019). The results are measured by accuracy with the reference answer.
  • Closed-domain QA on BoolQ(Clark et al. 2019). The results are measured by exact match.
  • Dialogue on MuTual(Cui et al. 2020). The results are measured by selecting one best-matched response from four available choices.
  • Summarization on SAMSum (Cui et al. 2020). The results are measured by the average of ROUGE-1, ROUGE-2, and ROGUE-L.
  • Named Entity Recognition on CoNLL03(Sang and De Meulder 2003). The results are measured by entity-level F1-score.
  • Sentiment Analysis on SST2(Socher et al. 2013). Results are measured by the accuracy of two-way classification.
Inference Efficiency
We demonstrate the inference efficiency of ELDER by reporting the inference latency. Our results show that ELDER maintains a similar inference speed to the base model, indicating its effectiveness for practical usage. The statistics are measured after 1000 edits from the ZsRE dataset.
GPT2-XL LLaMA2-7B
Base 0.56 s/edit 1.01 s/edit
ELDER 0.58 s/edit 1.04 s/edit
Table 6: Inference Efficiency
Detailed Results of Test Retention on General Tasks
The detailed results of the pre-edit and post-edit models on downstream tasks under each experimental setting are shown in Table 7, Table 8, Table 9 and Table 10. Each task has been described before. The last column in each table stands for the standard Test Retention on General Tasks we report in the main paper.
@@
Table 7: Results on general task of LLaMA2-7B after editing with ZsRE